Find the amount of ice cream in a cone if the radius of the cone is 4cm and its height is 10cm. The ice cream fully fills the cone and the hemisphere of the ice cream on top has a radius of 4 cm
Answer by edjones(8007)

V(cone)=pi*r^2*h/3
V(sphere)=(4/3)*pi*r^3 a hemisphere = 1/2 sphere
.
pi*4^2*10/3
=pi*16*10/3
=pi*160/3
167.552
.
4/6*pi*4^3 we use 4/6 instead of 4/3 because it is 1/2 a sphere
=(4/6)*pi*64
=(128/3)*pi
=134.041
.
167.552+134.041=301.593 cc
.
Ed